# Lead Times in Flux: Analyzing Airbnb Dynamics During Global Upheavals (2018-2022)

### v1

@@ 1 Introduction
Prior research has underscored the importance of lead times in tourism management. Lim and McAleer (2002) showed how lead times affect revenue management strategies in the hospitality sector. Tussyadiah and Pesonen (2016) examined consumer decision-making processes in tourism, finding that booking lead times are influenced by a complex interplay of factors, including perceived risk, travel purpose, and technological advancements. Lead times are frequently incorporated into forecasts of aggregate demand (Fiori and Foroni, 2019; Pereira, 2016; Huang and Zheng, 2023; Zaki, 2022; Fiori and Foroni, 2020; Webb et al., 2022), while Katz et al. (2024) and de Oliveira et al. (2021) presented models for forecasting lead times directly.
Global events, particularly crises such as pandemics, have profound effects on travel behaviors and lead times. The COVID-19 pandemic, in particular, has led to unprecedented disruptions in the tourism industry (Sainaghi and Chica-Olmo, 2022; Sharma et al., 2021; Qiu et al., 2020; Mueller and Sobreira, 2024; Okafor et al., 2022). According to Gössling et al. (2021), the pandemic resulted in massive declines in international and domestic travel, altering consumer risk perceptions and booking behaviors. Sigala (2020) emphasized that COVID-19 not only impacted demand but also accelerated digital transformation and the adoption of flexible booking policies in the industry. Research suggests that travelers’ perceived risk and uncertainty lead to significant changes in booking patterns. For example, Lo et al. (2011) found that during health crises, travelers tend to delay bookings or opt for destinations perceived as safer. Fuchs and Reichel (2011) showed that risk reduction strategies, such as flexible cancellation policies and enhanced safety measures, become more critical in influencing booking lead times.
Despite recognition of these shifts, much of the existing literature relies on traditional statistical measures—such as mean and median lead times—to analyze booking behaviors. While these measures are useful, they may not fully capture the distributional changes and increased variability introduced by external shocks like pandemics. As Assaf and Tsionas (2019a, b, c) argued, there is a need for more advanced analytical approaches in tourism research to better understand complex consumer behaviors.
@@
The primary objective of this research is to assess the effectiveness of the normalized L1 distance metric in detecting and quantifying changes in Airbnb booking lead times during periods of global upheaval, specifically the COVID-19 pandemic. By analyzing data from Airbnb properties across four major U.S. urban centers—Austin, Boston, Miami, and San Francisco—from 2018 to 2022, we aim to provide a comprehensive perspective on the evolving dynamics of urban tourism.
Furthermore, this study contributes to the literature by offering empirical evidence on how external shocks impact booking behaviors in the sharing economy context. By situating our findings within the broader theoretical frameworks of risk perception and consumer behavior during crises, we enhance the understanding of how travelers adjust their planning in response to heightened uncertainty.
The normalized L1 distance is used to quantify the divergence in lead time distributions across different years, offering a non-parametric measure that adjusts for scale differences by normalizing the traditional L1 (Manhattan) distance. This approach gives a global summary of lead time distributional changes, allowing a more nuanced understanding of how booking patterns have shifted in response to significant events. This framework provides practical insights for industry stakeholders and also advances methodological approaches in tourism research.
@@ 2.1 Data and Variable of Interest
We analyze Airbnb booking data for four major U.S. cities (Austin, Boston, Miami, and San Francisco) from January 2018 through December 2022. For each city, we consider both destination bookings (nights stayed in the city) and origin bookings (nights booked from the city), distinguishing between domestic and international travel corridors.
@@
Define:
[…]
  • $\Delta\in\{0,1,\ldots,365\}$: the lead time in days between the booking date and the check-in date.
  • $\ell_{i}$: the number of nights in booking $i$.
  • $\Delta_{i}$: the lead time (in days) for booking $i$.
We first collect all bookings that start in month $t$. For each booking $i$, we note its lead time $\Delta_{i}$ and length of stay $\ell_{i}$. Then, for each $\Delta$,
$w_{t}(\Delta)\;=\;\sum_{\begin{subarray}{c}i:\\ \Delta_{i}=\Delta\end{subarray}}\ell_{i},$
is the total number of nights that began in month $t$ with a lead time of $\Delta$ days. [^1] In other words, if two guests each book 5-night stays with $\Delta_{i}=3$, that adds $5+5=10$ to $w_{t}(3)$. Since $\Delta$ ranges from 0 to 365, this yields $366$ lead-time categories. To express these counts as a probability distribution, we normalize:
@@
Thus, $L_{t}(\Delta)$ is the proportion of all nights that start in month $t$ which were booked $\Delta$ days in advance. By construction,
@@ 2.3 Divergence Analysis Using Normalized L1 Distance
$D_{\text{L1}}(p,q)\;=\;\sum_{\Delta=0}^{365}\bigl{|}p(\Delta)-q(\Delta)\bigr{|}.$
@@
$D(p,q)\;=\;\frac{1}{2}\sum_{\Delta=0}^{365}\bigl{|}p(\Delta)-q(\Delta)\bigr{|}.$
@@
    $l^{1}_{{\text{YoY}},t}\;=\;\frac{1}{2}\sum_{\Delta=0}^{\Delta_{\max}}\Bigl{|}L_{t}(\Delta)-L_{t-12}(\Delta)\Bigr{|}.$
@@
    $l^{1}_{{\text{YoY2018}},t}\;=\;\frac{1}{2}\sum_{\Delta=0}^{\Delta_{\max}}\Bigl{|}L_{t}(\Delta)-L_{2018,t}(\Delta)\Bigr{|},$
@@ 4.1 Evolution of Lead Times
The analysis of Airbnb booking data from 2018 to 2022 shows significant variations in lead times across the cities of Austin, Boston, Miami, and San Francisco. We examined the mean, median, and standard deviation (SD) of lead times, segmented by domestic and international travel, and categorized by origin (bookings from the city) and destination (bookings to the city). Figures 3 and 4 show the trends in mean and median lead times, while tables 2 and 3 provide detailed yearly summaries.
@@ 4.2 Divergence Analysis Using Normalized L1 Distance
To quantify the changes in booking lead time distributions over time, we used the normalized L1 distance metric, calculating both the divergence from the baseline year 2018 ($l^{1}_{{\text{YoY2018}},t}$) and the year-over-year divergence ($l^{1}_{{\text{YoY}},t}$). Figures 5-8 display these divergences, with detailed statistics summarized in Tables 8-11.
@@ 4.2.1 Baseline Comparison Divergence ($l^{1}_{{\text{YoY2018}},t}$)
The divergence from the baseline year 2018 shows significant shifts in booking behaviors due to the pandemic. For international travel, divergence peaked in 2020 for all cities, indicating substantial disruptions. For example, the median $l^{1}_{{\text{YoY2018}},t}$ value for international bookings to Boston increased from 0.05 in 2019 to 0.34 in 2020. Similarly, for bookings originating from Austin, the median divergence increased from 0.05 in 2019 to 0.17 in 2020. Although there was some reduction in divergence in 2021 and 2022, the values remained higher than pre-pandemic levels, suggesting continued adjustments in international travel behaviors.
For domestic travel, divergence from the 2018 baseline increased in 2020 and remained elevated through 2022. For instance, the median $l^{1}_{{\text{YoY2018}},t}$ for domestic bookings from San Francisco increased from 0.03 in 2019 to 0.11 in 2020 and remained around 0.11–0.12 in subsequent years. In Boston, the divergence for domestic destination bookings increased from 0.03 in 2019 to 0.10 in 2020, rising further to 0.15 in 2022. These sustained divergences indicate prolonged changes in domestic booking behaviors compared to the pre-pandemic period.
@@ 4.2.2 Year-over-Year Divergence ($l^{1}_{{\text{YoY}},t}$)
The year-over-year divergence analysis shows significant temporal shifts in booking patterns. For domestic travel, $l^{1}_{{\text{YoY}},t}$ values peaked in 2021 for all cities, reflecting ongoing adjustments even as travel began to recover. For example, the median $l^{1}_{{\text{YoY}},t}$ for domestic bookings to San Francisco increased from 0.04 in 2019 to 0.12 in 2020 and further to 0.19 in 2021. In Austin, the median $l^{1}_{{\text{YoY}},t}$ for domestic bookings from the city increased from 0.04 in 2019 to 0.09 in 2020 and 0.13 in 2021.
For international travel, divergences peaked in either 2020 or 2021. The median $l^{1}_{{\text{YoY}},t}$ for international bookings from Miami increased from 0.04 in 2019 to 0.13 in 2020, with a slight decrease to 0.11 in 2021. For bookings to Boston, the median divergence increased from 0.05 in 2019 to 0.33 in 2020 and 0.39 in 2021. By 2022, divergences began to decrease across all cities, indicating a gradual stabilization of booking behaviors.
These patterns underscore the disruptions caused by the COVID-19 pandemic and the varied recovery trajectories across different cities and travel types. The elevated divergences in 2021 compared to 2020 for domestic travel show a ”two-phase” anomaly phenomenon, where the anomalous year 2020 serves as a new baseline, leading to higher $l^{1}_{{\text{YoY}},t}$ values even as travel patterns start to stabilize.
@@ 4.2.3 Correlation Analysis of Lead Time Divergences
The correlations between domestic and international $l^{1}$ divergences across cities and over time are shown in figures 9 and 10. High correlations within the same travel type and direction suggest consistent behavioral shifts, while lower correlations between different travel types indicate distinct factors influencing domestic and international booking patterns.
In Miami, there was a strong correlation between domestic and international divergences, suggesting that local conditions similarly affected both travel segments. For instance, Miami’s domestic origin $l^{1}_{{\text{YoY2018}},t}$ had a correlation of 0.83 with its international destination divergence. In contrast, San Francisco exhibited lower correlations between domestic and international travel, implying that different factors influenced booking behaviors in these segments.
@@ 4.2.4 Seasonal-Trend Decomposition (STL) Analysis
The Seasonal-Trend decomposition using LOESS (STL) on the normalized L1 distance time series for both the baseline year 2018 ($l^{1}_{{\text{YoY2018}},t}$) and the year-over-year ($l^{1}_{{\text{YoY}},t}$) divergences are shown in figures 1114. The trend components from the STL give insights into the persistent changes in booking lead time distributions over time.
For the $l^{1}_{{\text{YoY2018}},t}$ divergences, the trend components for both domestic and international travel peaked around mid-2020 to mid-2021, corresponding with the peak impact of the COVID-19 pandemic on global travel. These trend components do not return to zero by 2022, indicating that booking lead time distributions have not reverted to their pre-pandemic (2018) patterns. For example, the trend components for domestic origin divergences in Boston and San Francisco remain elevated going into 2023, suggesting sustained shifts in booking behaviors in these markets. This persistent divergence implies that travelers continue to book accommodations with lead times significantly different from those in 2018, reflecting lasting changes in planning behavior possibly due to altered risk perceptions, remote work arrangements, or new travel preferences.
In contrast, the $l^{1}_{{\text{YoY}},t}$ divergences show a different pattern. For domestic travel, the trend components generally decrease towards pre-pandemic levels by 2022, indicating that year-over-year changes in booking behaviors are stabilizing. However, the divergences remain higher than zero, suggesting that while the rate of change is normalizing, booking patterns have not fully returned to previous norms. For international origin travel, the trend components show an acceleration of divergence going into 2023. This indicates that international booking behaviors continue to change significantly compared to the immediate past year.
These findings show that booking lead time distributions have undergone significant and lasting changes since the onset of the pandemic. These changes are not uniformly reverting to pre-pandemic patterns, especially in international travel, showing the importance of using comprehensive divergence metrics to inform industry responses and strategies.

### v3

@@ 1 Introduction
Prior research has underscored the importance of lead times in tourism management. Lim and McAleer (2002) showed how lead times affect revenue management strategies in the hospitality sector. Tussyadiah and Pesonen (2016) examined consumer decision-making processes in tourism, finding that booking lead times are influenced by a complex interplay of factors, including perceived risk, travel purpose, and technological advancements. Lead times are frequently incorporated into forecasts of aggregate demand (Fiori and Foroni, 2019; Pereira, 2016; Zaki, 2022; Fiori and Foroni, 2020; Webb et al., 2022), while Katz et al. (2024) and de Oliveira et al. (2021) presented models for forecasting lead times directly. Moreover, Song and Li (2008) and Huang and Zheng (2023) have reviewed the broader context of tourism demand forecasting, emphasizing the growing importance of incorporating booking windows into predictive frameworks.
Global events, particularly crises such as pandemics, have profound effects on travel behaviors and lead times. The COVID-19 pandemic, in particular, has led to unprecedented disruptions in the tourism industry (Sainaghi and Chica-Olmo, 2022; Sharma et al., 2021; Qiu et al., 2020; Mueller and Sobreira, 2024; Okafor et al., 2022). According to Gössling et al. (2021), the pandemic resulted in massive declines in international and domestic travel, altering consumer risk perceptions and booking behaviors. Sigala (2020a) emphasized that COVID-19 not only impacted demand but also accelerated digital transformation and the adoption of flexible booking policies in the industry. Research suggests that travelers’ perceived risk and uncertainty lead to significant changes in booking patterns. For example, Lo et al. (2011) found that during health crises, travelers tend to delay bookings or opt for destinations perceived as safer. Fuchs and Reichel (2011) showed that risk reduction strategies, such as flexible cancellation policies and enhanced safety measures, become more critical in influencing booking lead times.
Additionally, recent research in the airline industry underscores similar shifts in booking lead-time patterns under COVID-19. Zhang et al. (2021) investigated the Chinese aviation market and revealed how pandemic-related uncertainties drove travelers to shorten planning horizons and switch travel modes. Their moment-based analysis demonstrates that variance and higher-order distributional features of lead times can provide actionable insights into passenger behavior, a perspective that aligns well with the comprehensive approach advocated in this paper.
Despite recognition of these shifts, much of the existing literature relies on traditional statistical measures—such as mean and median lead times—to analyze booking behaviors. While these measures are useful, they may not fully capture the distributional changes and increased variability introduced by external shocks like pandemics. As Assaf and Tsionas (2019a, b, d) argued, there is a need for more advanced analytical approaches in tourism research to better understand complex consumer behaviors. Likewise, Assaf and Tsionas (2019c) underscored the importance of distribution-aware metrics in examining tourism demand, especially during times of heightened uncertainty.
@@
The primary objective of this research is to assess the effectiveness of the normalized L1 distance metric in detecting and quantifying changes in Airbnb booking lead times during periods of global upheaval, specifically the COVID-19 pandemic. By analyzing data from Airbnb properties across four major U.S. urban centers—Austin, Boston, Miami, and San Francisco—from 2018 to 2022, we aim to provide a comprehensive perspective on the evolving dynamics of urban tourism. Our work also resonates with the emerging body of research on digital platforms’ responses to crises; as noted by Tussyadiah and Pesonen (2016) and Gyódi (2021), peer-to-peer accommodation providers (like Airbnb) can exhibit different booking behaviors compared to traditional hotels, a gap that demands further empirical and methodological attention.
Furthermore, this study contributes to the literature by offering empirical evidence on how external shocks impact booking behaviors in the sharing economy context. By situating our findings within the broader theoretical frameworks of risk perception and consumer behavior during crises, we enhance the understanding of how travelers adjust their planning in response to heightened uncertainty. Our focus on Airbnb properties answers recent calls (e.g., Sigala, 2020b) to examine how platform-specific features—such as flexible listing types, decentralized inventory, and digital-first user experiences—shape travel demand under extreme market disruptions.
The normalized L1 distance is used to quantify the divergence in lead time distributions across different years, offering a non-parametric measure that adjusts for scale differences by normalizing the traditional L1 (Manhattan) distance. This approach gives a global summary of lead time distributional changes, allowing a more nuanced understanding of how booking patterns have shifted in response to significant events. This framework provides practical insights for industry stakeholders and also advances methodological approaches in tourism research, as it addresses the need for enhanced analytics that capture full distributional variations rather than focusing on singular statistical descriptors alone.
@@ 2.1 Data and Variable of Interest
We analyze Airbnb booking data for four major U.S. cities (Austin, Boston, Miami, and San Francisco) from January 2018 through December 2022. For each city, we consider both destination bookings (nights stayed in the city) and origin bookings (nights booked from the city), distinguishing between domestic and international travel corridors. In total, the dataset comprises monthly aggregates of booking records, where each record indicates how many nights were booked in a given city (or from a given city) for specific calendar months of stay.
A key variable in our study is the lead time, defined as the number of days between the booking date and the check-in date. We confine our analysis to lead times from 0 to 365 days before check-in. Although travelers can book more than one year in advance, these cases represent under .5% of total bookings in our dataset and are thus negligible. By restricting the lead-time domain to 0–365 days, we ensure that more than 99% of all nights are captured, while also simplifying the analysis (e.g., a uniform bin structure for the lead-time distribution).
@@
Let $B_{t}$ be the set of all bookings whose check-in date (i.e., first night) falls in month $t$. For each booking $i\in B_{t}$,
  • $\Delta_{i}\in\{0,1,\ldots,365\}$ denotes the lead time (in days) between its booking date and its check-in date,
  • $\ell_{i}$ denotes the number of nights in booking $i$.
For each lead time $\Delta\in\{0,\dots,365\}$, define
$w_{t}(\Delta)\;=\;\sum_{\begin{subarray}{c}i\in B_{t}\\ \Delta_{i}=\Delta\end{subarray}}\ell_{i},$
that is, the total number of nights contributed by all bookings in $B_{t}$ whose lead time is exactly $\Delta$ [^1] For example, if two guests each book 5-night stays with $\Delta_{i}=3$, then they add $5+5=10$ to $w_{t}(3)$.
Since $\Delta$ ranges from 0 to 365, we have $366$ possible lead-time categories. To convert these totals into a probability distribution, we normalize:
@@
Hence, $L_{t}(\Delta)$ represents the proportion of all nights (among bookings that start in month $t$) that were booked $\Delta$ days in advance. By construction,
@@ 2.3 Divergence Analysis Using Normalized L1 Distance
$D_{\text{L1}}(p,q)\;=\;\sum_{\Delta=0}^{365}|p(\Delta)-q(\Delta)|.$
@@
$D(p,q)\;=\;\frac{1}{2}\sum_{\Delta=0}^{365}|p(\Delta)-q(\Delta)|.$
@@
Figure 1: Example of two lead-time distributions, A (left) and B (right), that share the same mean ($\mu=188$) and standard deviation ($\sigma=30$), but exhibit notably different shapes. Distribution A is roughly unimodal, while Distribution B is bimodal. Despite having identical mean and SD, the two distributions have an L1 distance of $0.25$, highlighting how summary statistics alone may fail to capture important shifts.
Figure 1 illustrates why L1 distance can be more informative. Distributions A and B both have mean $188$ and standard deviation $30$, yet differ substantially in shape (unimodal vs. bimodal). Commonly used summary statistics do not reveal these differences, whereas the normalized L1 distance reveals a notable shift: here, $D(A,B)=0.25$, indicating that 25% of the probability mass has effectively “moved.”
@@
    $l^{1}_{{\text{YoY}},t}\;=\;\frac{1}{2}\sum_{\Delta=0}^{\Delta_{\max}}|L_{t}(\Delta)-L_{t-12}(\Delta)|.$
@@
    $l^{1}_{{\text{YoY2018}},t}\;=\;\frac{1}{2}\sum_{\Delta=0}^{\Delta_{\max}}|L_{t}(\Delta)-L_{2018,t}(\Delta)|,$
@@ 3.1 Setup
Figure 2: The distribution of lead times for the fictional town B-Ville, illustrating the changes in booking behaviors across two consecutive years. This figure shows how the normalized L1 distance metric captures small perturbations in lead times better than traditional summary statistics like mean, median, or standard deviation, thereby emphasizing its sensitivity to distributional changes in booking behaviors.
@@ 3.2 Results and Effects on the Pickup Method
Figure 3: Pickup curves (cumulative sum of lead time distributions) for B-Ville in 2019 and 2020. The 2020 pickup curve indicates a slower accumulation of bookings compared to 2019, leading to forecasting errors when using the pickup method.
@@
Year 1 Year 2 Change (%)
Mean 12.38066 12.82447 3.58
Median 8 8 0
Standard Deviation 9.417028 9.012016 -4.30
L1 Distance Change 0.2156462
Table 1: Results from simulated lead times data in B-Ville. The L1 distance change better captures the distributional shifts compared to mean, median, or standard deviation alone.
@@ 4 Airbnb Results
In this section, we present both conventional summary metrics (mean, median, standard deviation) and the normalized L1 divergence for booking lead times across four U.S. cities (Austin, Boston, Miami, and San Francisco) from 2018 to 2022. Our aim is to show how these two approaches complement each other in revealing shifts in booking behavior—while mean or median changes capture overall trends in travelers’ planning horizons, the L1 metric uncovers subtle distributional movements that may not be reflected in central tendencies.
@@ 4.1 Evolution of Lead Times
The analysis of Airbnb booking data from 2018 to 2022 shows significant variations in lead times across the cities of Austin, Boston, Miami, and San Francisco. We examined the mean, median, and standard deviation (SD) of lead times, segmented by domestic and international travel, and categorized by origin (bookings from the city) and destination (bookings to the city). Figures 4 and 5 show the trends in mean and median lead times, while Supplementary Tables S1 and S2 provide detailed yearly summaries.
@@
Figure 4: Trends in mean (solid line) and median (dashed line) lead times for Airbnb bookings to Austin, Boston, Miami, and San Francisco from 2018 to 2022. The figure differentiates domestic (red) and international (blue) travel, showing how booking behaviors evolved over time, particularly during significant events such as the COVID-19 pandemic.
Figure 5: Trends in mean (solid line) and median (dashed line) lead times for Airbnb bookings from Austin, Boston, Miami, and San Francisco from 2018 to 2022. The figure differentiates domestic (red) and international (blue) travel, showing how booking behaviors evolved over time, particularly during significant events such as the COVID-19 pandemic.
@@ 4.2 Divergence Analysis Using Normalized L1 Distance
To capture shifts across the entire lead-time distribution, we use the normalized L1 distance, calculated both relative to the baseline year 2018 ($l^{1}_{{\text{YoY2018}},t}$) and relative to the same month in the previous year ($l^{1}_{{\text{YoY}},t}$). Figures 67 depict the baseline comparisons, while Figures 89 illustrate the year-over-year divergences. The associated numeric summaries appear in Supplementary Tables S7S10.
@@ 4.2.1 Baseline Comparison Divergence ($l^{1}_{{\text{YoY2018}},t}$)
Comparisons with 2018 underscore how dramatically the pandemic reshaped booking behaviors. For international travel, $l^{1}_{{\text{YoY2018}},t}$ peaked in 2020 across all four cities, indicating substantial distributional disruption. In Boston, for instance, the median divergence for international destination bookings rose from 0.05 in 2019 to 0.34 in 2020, while in Austin it climbed from 0.05 to 0.17. Although these divergences eased somewhat in 2021 and 2022, they generally remained elevated compared to pre-pandemic levels, suggesting that international booking patterns never fully reverted to their 2018 norms.
A similar story emerges for domestic bookings: San Francisco’s median $l^{1}_{{\text{YoY2018}},t}$ increased from 0.03 in 2019 to 0.11 in 2020 and stabilized around 0.11–0.12 thereafter, while Boston’s divergence rose from 0.03 to 0.10 and ultimately reached 0.15 by 2022. These sustained gaps indicate that, even when the mean or median domestic lead time did not exhibit a massive year-to-year swing, parts of the distribution migrated toward notably shorter or longer booking windows. As a result, the overall distribution in 2022 still looked markedly different from its 2018 baseline.
Figure 6: Normalized L1 divergence in destination lead times for Airbnb bookings, taking 2018 as the baseline year.
Figure 7: Normalized L1 divergence in origin lead times for Airbnb bookings, taking 2018 as the baseline year.
@@ 4.2.2 Year-over-Year Divergence ($l^{1}_{{\text{YoY}},t}$)
Year-over-year divergences (Figures 89) highlight temporal shifts within the pandemic period. For domestic travel (red lines), $l^{1}_{{\text{YoY}},t}$ often surged in 2021 as the industry began recovering, then tapered in 2022. In San Francisco, the median $l^{1}_{{\text{YoY}},t}$ for domestic destination bookings jumped from 0.04 in 2019 to 0.12 in 2020 and 0.19 in 2021, mirroring the transitions seen in the simpler metrics but adding clarity on how much of the distribution had changed. Meanwhile, in Miami, the domestic origin year-over-year divergence rose from 0.04 in 2019 to 0.09 in 2020 and 0.13 in 2021, underscoring ongoing flux.
For international bookings (blue lines), peaks occurred in either 2020 or 2021. The median $l^{1}_{{\text{YoY}},t}$ for bookings from Miami, for instance, reached 0.13 in 2020 before easing slightly; in Boston, the international destination divergence swelled from 0.05 in 2019 to 0.33 in 2020 and 0.39 in 2021. Although these values generally drifted downward in 2022, they remained well above pre-pandemic norms, reflecting prolonged turbulence in international travel. Indeed, the “two-phase” anomaly—where 2020 itself becomes a new, unconventional baseline—left many city pairs experiencing persistently high year-over-year divergences in 2021.
Figure 8: Normalized L1 divergence in destination lead times for Airbnb bookings, taking the previous year as the baseline year.
Figure 9: Normalized L1 divergence in origin lead times for Airbnb bookings, taking the previous year as the baseline year.
@@ 4.2.3 Correlation Analysis of Lead Time Divergences
The correlations between domestic and international $l^{1}$ divergences across cities and over time are shown in Supplementary Figures S1 and S2. High correlations within the same travel type and direction suggest consistent behavioral shifts, while lower correlations between different travel types indicate distinct factors influencing domestic and international booking patterns.
In Miami, there was a strong correlation between domestic and international divergences, suggesting that local conditions similarly affected both travel segments. For instance, Miami’s domestic origin $l^{1}_{{\text{YoY2018}},t}$ (bottom-left panel of Supplementary figure S1) had a correlation of 0.83 with its international destination divergence. In contrast, San Francisco (bottom-right panel of Supplementary figure S1) exhibited lower correlations between domestic and international travel, implying that different factors influenced booking behaviors in these segments.
@@ 4.2.4 Seasonal-Trend Decomposition (STL) Analysis
The Seasonal-Trend decomposition using LOESS (STL) on the normalized L1 distance time series for both the baseline year 2018 ($l^{1}_{{\text{YoY2018}},t}$) and the year-over-year ($l^{1}_{{\text{YoY}},t}$) divergences are shown in Supplementary Figures S3S6. The trend components from the STL give insights into the persistent changes in booking lead time distributions over time.
For the $l^{1}_{{\text{YoY2018}},t}$ divergences (Supplementary Figures S3-S4), the trend components for both domestic and international travel peaked around mid-2020 to mid-2021, corresponding with the peak impact of the COVID-19 pandemic on global travel. These trend components do not return to zero by 2022, indicating that booking lead time distributions have not reverted to their pre-pandemic (2018) patterns. For example, the trend components for domestic origin divergences in Boston and San Francisco remain elevated going into 2023, suggesting sustained shifts in booking behaviors in these markets. This persistent divergence implies that travelers continue to book accommodations with lead times significantly different from those in 2018, reflecting lasting changes in planning behavior possibly due to altered risk perceptions, remote work arrangements, or new travel preferences.
In contrast, the $l^{1}_{{\text{YoY}},t}$ divergences (Supplementary Figures S5-S6) show a different pattern. For domestic travel, the trend components generally decrease towards pre-pandemic levels by 2022, indicating that year-over-year changes in booking behaviors are stabilizing. However, the divergences remain higher than zero, suggesting that while the rate of change is normalizing, booking patterns have not fully returned to previous norms. For international origin travel, the trend components show an acceleration of divergence going into 2023. This indicates that international booking behaviors continue to change significantly compared to the immediate past year.